OpenStudy (loser66):

Show that if n is a composite integers with n is not 4, then (n-1) ! =0 mod n Please, help

OpenStudy (loser66):

I don't know how to do it but: n! = n (n-1)! = 0 mod n n = a. b for some a, b in N, and \(a, b\cancel \equiv 0\) hence only one way to get it hold is \((n-1)! \equiv 0 (mod n)\) Am I crazy?

OpenStudy (superdavesuper):

dat should work. another way to prove is that n=a.b for some a,b in N and a,b<n so (n-1)! = 1...a..b..(n-1) for n>4 then (n-1)! = 0 mod n will follow

OpenStudy (loser66):

Nope, I don't think so a.b = n , hence a.b > n-1 both them cannot be on (n-1) !

OpenStudy (superdavesuper):

hahahaa sorry - its a comma , not period . alright a, b in N and n>a, n>b so for n>4, (n-1)!=1...a...b..(n-1) happy now? :)

OpenStudy (loser66):

no :)

OpenStudy (loser66):

\(n = a*b\rightarrow a < n\rightarrow a \leq n-1\rightarrow a| (n-1)!\) Same to b \(n = a*b\rightarrow b < n\rightarrow b \leq n-1\rightarrow b| (n-1)!\) But nothing to guarantee that a.b | (n-1)! like 4 | 8 8|8 but 32 not | 8

OpenStudy (superdavesuper):

ahhhh....so difficult to put this in writing....im not saying a.b < n or a.b < (n-1) u are right: it is NOT but cuz' n=a.b and n>a and n>b AND a and b are not equal to (n-1) (something about consecutive numbers cannot have a common factor except 1) SO n-1>a and n-1>b it follows that (n-1)!=1...a..b..(n-1)

OpenStudy (loser66):

Let try other way: Let \( n = p1^{a_1} p2^{a_2}........pr^{a_r}\) each of pi < n , hence \(p_i^{a_i} \leq n-1\) Hence by Euclid, \((n-1)! \equiv 0 (mod ~~p_i^{a_i}\)) And \((p_i^{a_i}, p_j ^{a_j}) =1\) Hence by Wilson's theorem \((n-1)! \equiv 0 (mod~~lcm ~~(p_i^{a_i}))\) And the last notation is n

OpenStudy (superdavesuper):

both of ur proves should work :) but the thing that bugs me is that somewhere there should be a restriction that n is not equal to 4 as given.

OpenStudy (loser66):

If n = 4, then only one composition is 2*2 , then n -1 = 3 and 3! =6 and then 6 not | 4, so that \(6 \cancel \equiv 0 (mod 4)\)

OpenStudy (superdavesuper):

just weird that its not intrinsic to ur proves but they will work :)